OpenStudy (anonymous):

7^2x=5^5+8y. how'd i solve for y? Please help

OpenStudy (anonymous):

So, \[7^{2}x=5^{5}+8y\]

OpenStudy (anonymous):

yep.. so how'd i solve for Y?

OpenStudy (anonymous):

Well first you would do exponents so, 7*7= 49 & 5*5*5*5*5 = 3125 so, \[49x=3125+8y\]

OpenStudy (anonymous):

That's as much as you can do, I need to know what x or y is to finish the equation it is physically impossible to solve the equation.

OpenStudy (anonymous):

okay then.. Thank you!